Please help me with my second to last problem of my homework! I really need help...
To go over a tree, a golfer uses a 7-iron to hit the ball with an initial upward velocity of 80 feet per second.
a. Write a quadratic function that gives the height of the ball t seconds after it is hit.
I think that a. is h(t)= 80t-16t^2 but I'm not sure...
b. Graph the function you found in part (a). (you can only give me the T chart.)
c. After how many seconds will the ball be at its maximum height?
d. What is the approximate maximum height the ball will reach?

Answer by nerdybill(7384)

To go over a tree, a golfer uses a 7-iron to hit the ball with an initial upward velocity of 80 feet per second.
a. Write a quadratic function that gives the height of the ball t seconds after it is hit.
I think that a. is h(t)= 80t-16t^2 but I'm not sure... (YES!)
b. Graph the function you found in part (a). (you can only give me the T chart.)
c. After how many seconds will the ball be at its maximum height?
t = -b/(2a)
t = -80/(2(-16))
t = -80/(-32)
t = 2.5 seconds
d. What is the approximate maximum height the ball will reach?
plug t=2.5 from b into original equation:
h(2.5)= 80(2.5)-16(2.5)^2
h(2.5)= 98.4 feet